# GaussianEM: Model compositional and conformational heterogeneity using 3D Gaussians

### v1

@@ Results
After training, we performed PCA on the latent space to explore conformational variability. The first principal component revealed discrete conformational transitions, as illustrated in Figure 5b. Three representative volumes were reconstructed, two of which correspond to known structures. The gray volume represents the main body of the VgrG spike, consistent with the atomic model PDB-9JG9. The yellow conformation includes the BtpeA and BtaeB effectors assembled on the base of the gray structure, corresponding to PDB-9JG8. The blue conformation further extends the state, maintaining the additional structural region mentioned above and highlighted by the red box. To validate these results, the three clustered projections were reconstructed in CryoSPARC (Figure 5c). Although the additional region remains partially unresolved, the reconstructions reveal clearer envelope features around the effectors, as indicated by the dashed red box. Moreover, the second and third principal components in Figure 5d capture continuous motions within the effector region. This change was easily neglected (see Figure S3 in Supplementary Material section “T6SS effectors”). The reconstructed volumes along these components reveal smooth structural transitions rather than discrete states. Gaussian displacements were mapped onto the assembled atomic models, where stable regions are colored gray and flexible regions are shown in red and blue for visualization. Interestingly, the reconstructed volumes along the second component differ from those along the third: the former retain additional peripheral structures, while the latter preserve the core architecture of the two effectors and the VgrG spike.
@@
The $\alpha V\beta 8$ integrin is a highly flexible protein involved in cell differentiation during development (EMPIAR-10345) campbell2020cryo. This dataset consists of 84266 particle images with a box size of 300 pixels of width 1.35 Å. The flexible arm of the integrin is difficult to resolve through conventional homogeneous reconstruction; therefore, we used the density map refined by 3DFlex punjani20233dflex as our initial model. All Gaussian parameters were allowed to vary during training. Using K-means clustering, five representative volumes were identified and reconstructed in Figure 6b. In addition to the arm-swing motion previously reported punjani20233dflex, GaussianEM also revealed the absence of the arm and the foot of the core protein. These findings were validated by homogeneous reconstruction in CryoSPARC using clustered projections (Figure 6c). Furthermore, with the help of compact clustering, the motion of the arm is also visible in CryoSPARC reconstruction. Figure 6 illustrates the detailed conformational changes of $\alpha V\beta 8$ integrin along the first and third principal components.
Quantitative comparison on conformational motions
As the inaugural public challenge in the field, the Flatiron Institute (FI) Cryo-EM Conformational Heterogeneity Challenge astore2025inaugural introduced a novel set of evaluation metrics and successfully attracted wide participation from leading research groups. To quantitatively compare the capabilities of different structural heterogeneity analysis methods in characterizing conformational motions, we adopt one of its key metrics, Percentage of Captured Variance (PCV), to calculate the overlap between the conformational subspaces generated by two distinct methods. Specifically, considering the conformational set from method A as the reference set and method B as the test set, the mathematical expression of PCV(A, B) can be formulated as:
@@
where $\mathbf{V}_{A}$ and $\mathbf{S}_{A}$ denote the eigenvector matrix and the diagonal eigenvalue matrix of the reference set’s covariance matrix, respectively, $\mathbf{V}_{B}$ represents the eigenvector matrix of the test set’s covariance matrix, and $\|\cdot\|_{F}$ is Frobenius norm. Essentially, the PCV metric quantifies the projection of the weighted subspace of the reference set that projects onto the subspace spanned by the test set. This ratio indicates the proportion of the reference set’s conformational variance that is accounted for by the test set, with a value range of $[0,1]$. The key to its analytical ability lies in its asymmetry. For instance, the superiority of Method B over Method A is established through an asymmetric result in the pairwise PCV comparison. This occurs when $\text{PCV}(A,B)$ is higher than $\text{PCV}(B,A)$. Such a result indicates that the conformational space of Method B effectively captures the variance within that of Method A, while the conformational space of Method A fails to cover the unique conformations explored by Method B. This outcome demonstrates that Method B provides a more comprehensive and diverse conformational map set.
Method IgG-1D
CryoDRGN DynaMight 3DFlex GaussianEM
GT 0.5982 0.6605
Table 1: PCV result on IgG-1D dataset. ”–” indicates that the method fails to produce a valid result.
To evaluate the ability of GaussianEM in resolving conformational motions, quantitative comparisons were performed against three widely used methods: CryoDRGN, DynaMight, and 3DFlex. All methods were trained using their default settings. Using the ground-truth (GT) conformations as a reference, we first assessed the conformational diversity recovered from the IgG-1D dataset. For each method, we generated an ensemble of 100 conformations by applying K-means clustering to the latent vectors and decoding the class centroids into 3D volumes. As summarized in Table 1, GaussianEM produces a more diverse conformational ensemble than CryoDRGN. In contrast, DynaMight and 3DFlex fail to effectively resolve conformations in datasets with large-scale structural motions.
@@
Next, we extended our evaluation to a diverse set of publicly available EMPIAR datasets. For each method, the conformational space was represented by 50 reconstructed conformations. As shown in Table 2, the results on the EMPIAR dataset reveal clear differences in the ability of the methods to explore conformational space. Specifically, the comparison between GaussianEM and CryoDRGN (dark blue) indicates that, GaussianEM recovers a broader and more diverse conformational space. This space encompasses the major motions identified by CryoDRGN, whereas CryoDRGN fails to capture some of the unique conformations revealed by GaussianEM. As illustrated in Figure 4e, GaussianEM successfully captures the principal conformational changes occurring in the dynamic domains of the precatalytic spliceosome. Similarly, the comparisons in the light green and light blue regions demonstrate the superiority of CryoDRGN over DynaMight and of DynaMight over 3DFlex, respectively. Overall, these results establish a consistent performance ranking: GaussianEM >> CryoDRGN >> 3DFlex. A similar ranking was observed on the other EMPIAR dataset.
@@
Finally, to further assess robustness and generalizability, we performed PCV comparisons on two datasets from the FI cryo-EM heterogeneity challenge, both exhibiting continuous conformational heterogeneity of thyroglobulin (Table 3). The first is an experimental dataset containing 33,742 particle images, and the second is a simulated dataset of the same size, generated by projecting 3,861 distinct conformations obtained from molecular dynamics simulations. The pairwise PCV results consistently show that GaussianEM outperforms CryoDRGN in resolving conformational heterogeneity, while CryoDRGN in turn outperforms 3DFlex. Notably, DynaMight failed to produce plausible conformations on both datasets.
@@ Discussion
In this study, we introduce GaussianEM, a Gaussian-based framework designed to intuitively capture both compositional and conformational heterogeneity in cryo-EM datasets. Most recent cryo-EM heterogeneity analysis methods typically rely on voxel-based representations of macromolecular structures. Such representations allow for identification of compositional variations by comparing density between reconstructed volumes. However, they are often limited in describing continuous conformational motion. Tracking subtle structural changes using voxel grids usually requires manual interpretation and lacks smoothness. Additionally, this process sometimes introduces artifacts. In contrast, Gaussian-based models offer a more natural way to describe structural variability. In this formulation, changes in Gaussian density directly correspond to the presence or absence of components, while Gaussian displacements provide an intuitive visualization of continuous motion. This representation makes it easier to locate and interpret regions undergoing change. Moreover, the isotropic Gaussian function ensures consistent behavior in both two and three dimensions, linking 2D projections directly to 3D structural deformations. These properties form the foundation of GaussianEM in exploring molecular heterogeneity.
@@ Methods
We represent the protein density map with a group of learnable 3D Gaussians $\mathbb{G}^{3}=\{G^{3}_{i}\}_{i=1,\ldots,N}$ kerbl20233d. Each Gaussian $\mathbb{G}^{3}_{i}$ depicts a local Gaussian-shaped density distribution, i.e.,
@@
A good property of Gaussian distribution is that its integral along one coordinate axis (usually $z$ row in the Cryo-EM domain) in 3D yields a 2D Gaussian distribution zha2024r; bae2024per. The $\hat{\mathbf{x}}\in\mathbb{R}^{2}$, $\hat{\mathbf{p}}_{i}\in\mathbb{R}^{2}$ and $\hat{\mathbf{\Sigma}}_{i}\in\mathbb{R}^{2\times 2}$ of projected Gaussians $\mathbb{G}^{2}_{i}$ are easily obtained by directly discarding the specific rows and columns of their corresponding 3D versions. The rendered particle projection $I_{k}\in\mathbb{R}^{H\times W}$ can be formulated as
@@
where $\mathcal{R}$ denotes the complete rendering process, $C_{k}$ denotes the CTF modulator, $P$ is the simple projection operator along the $z$ row, and $T$ transforms the 3D Gaussians given pose $\phi_{k}$. For computational efficiency, $T$ actually calculates the inverse matrix of particle pose to align the viewing direction with the coordinate axis.
@@
Consistent with most existing work, the heterogeneous reconstruction network adopts an encoder–decoder architecture that, maps raw particle images to their possible deformed volumes. Within the framework of 3D Gaussians, the deformation of flexible structures is characterized by the change of Gaussian parameters, whereas the invariant parameters correspond to the canonical density map obtained from homogeneous reconstruction. Our heterogeneous reconstruction network consists of three main modules, two encoders and one decoder, all implemented using multilayer perceptrons with Relu activation functions (see Supplementary Material section “Network architecture”). First, the raw particle images are processed by an image encoder made up of three hidden layers, which transforms them into $i$-dimension latent variables. In neural radiance field-based methods, position encoding is introduced to facilitate the network determining particle poses and learning high-frequency features. Similarly, we design a Gaussian encoder to assign a unique identifier to each Gaussian, helping to distinguish and aggregate different Gaussians. This encoder takes the density, scaling, position, and rotation parameters of each Gaussian as input and produces a $j$-dimension latent embedding. This per-Gaussian strategy has no limit on the number of Gaussians, allowing the incorporation of given atomic models and yielding more accurate estimations.
@@
Geometric Constraints. Overfitting can easily occur in the absence of well-designed regularization, particularly when dealing with high-noise experimental images. In such cases, Gaussians tend to aggregate, resulting in low-resolution structures that contribute uniformly across inputs. Considering the inherent continuity of protein motions, smoothness and local rigidity are essentail properties to maintain. For this purpose, we impose two types of constraints on the estimated parameters: one for Gaussian embedding and the other just on the Gaussian displacement. For Gaussian embedding $\mathbf{j}_{n}$, given the average value of nearest point distance $d_{mean}$, we consider the neighboring Gaussians within a radius of $\mu_{1}=2.5\times d_{mean}$;
@@
where $\mathcal{N}(n,\mu)$ denotes the set of Guassians within distance $\mu$ from Gaussian $n$, and $d_{n,s}$ is the initial space distance between Gaussians $n$ and $s$. The weight $\omega_{n,s}$ controls the strength of constraints. The closer the two Gaussians are, the stronger the constraint is. This embedding constraint encourages adjacent Gaussians to udnergo similar parameter changes, thereby preserving local consistency in both compositional and conformational heterogeneity. For the displacements of Gaussian, we first enforce the local rigidity of deformations by constraining the relative distance between adjacent Gaussians.

### v2

@@ Results
After training, we performed PCA on the latent space to explore conformational variability. The first principal component revealed discrete conformational transitions, as illustrated in Figure 5b. Three representative volumes were reconstructed, two of which correspond to known structures. The gray volume represents the main body of the VgrG spike, consistent with the atomic model PDB-9JG9. The yellow conformation includes the BtpeA and BtaeB effectors assembled on the base of the gray structure, corresponding to PDB-9JG8. The blue conformation further extends the state, maintaining the additional structural region mentioned above and highlighted by the red box. To validate these results, the three clustered projections were reconstructed in CryoSPARC (Figure 5c). Although the additional region remains partially unresolved, the reconstructions reveal clearer envelope features around the effectors, as indicated by the dashed red box. Moreover, the second and third principal components in Figure 5d capture continuous motions within the effector region. This change was easily neglected (see Figure LABEL:fig:gao_drgn in Supplementary Material section “T6SS effectors”). The reconstructed volumes along these components reveal smooth structural transitions rather than discrete states. Gaussian displacements were mapped onto the assembled atomic models, where stable regions are colored gray and flexible regions are shown in red and blue for visualization. Interestingly, the reconstructed volumes along the second component differ from those along the third: the former retain additional peripheral structures, while the latter preserve the core architecture of the two effectors and the VgrG spike.
@@
The $\alpha V\beta 8$ integrin is a highly flexible protein involved in cell differentiation during development (EMPIAR-10345) Campbell et al. (2020). This dataset consists of 84266 particle images with a box size of 300 pixels of width 1.35 Å. The flexible arm of the integrin is difficult to resolve through conventional homogeneous reconstruction; therefore, we used the density map refined by 3DFlex Punjani and Fleet (2023) as our initial model. All Gaussian parameters were allowed to vary during training. Using K-means clustering, five representative volumes were identified and reconstructed in Figure 6b. In addition to the arm-swing motion previously reported Punjani and Fleet (2023), GaussianEM also revealed the absence of the arm and the foot of the core protein. These findings were validated by homogeneous reconstruction in CryoSPARC using clustered projections (Figure 6c). Furthermore, with the help of compact clustering, the motion of the arm is also visible in CryoSPARC reconstruction. Figure 6 illustrates the detailed conformational changes of $\alpha V\beta 8$ integrin along the first and third principal components.
Quantitative comparisons
Quantifying the performance of heterogeneity analysis methods remains a major challenge in computational structural biology. Homogeneous structure determination benefits from established benchmarks such as the Fourier Shell Correlation (FSC) curve. In contrast, heterogeneous reconstruction lacks universally accepted evaluation metrics due to the inherent continuity and complexity of the conformational space. To address this limitation, several quantitative measures have been proposed, ranging from conformational diversity to voxel-level consistency. Building on these metrics, we performed a comprehensive quantitative comparison of GaussianEM with representative heterogeneous reconstruction methods, including CryoDRGN Zhong et al. (2021), DynaMight Schwab et al. (2023), and 3DFlex Punjani and Fleet (2023).
Completeness of heterogeneity analysis
The Flatiron Institute (FI) Cryo-EM Conformational Heterogeneity Challenge Astore et al. (2025) attracted broad participation from leading research groups and introduced a novel set of evaluation metrics for conformational heterogeneity. We adopt one of its key metrics, Percentage of Captured Variance (PCV), to quantitatively assess the ability of different methods to characterize conformational motions. PCV is a pairwise comparison metric that measures the overlap between two extracted latent subspaces. Specifically, when the conformational set produced by method A is treated as the reference, and that of method B as the test set, the PCV value PCV(A, B) is defined as:
@@
where $\mathbf{V}_{A}$ and $\mathbf{S}_{A}$ denote the eigenvector matrix and the diagonal eigenvalue matrix of the reference set’s covariance matrix, respectively, $\mathbf{V}_{B}$ represents the eigenvector matrix of the test set’s covariance matrix, and $\|\cdot\|_{F}$ is Frobenius norm. The PCV metric measures how much of the weighted conformational subspace of a reference set can be projected onto the subspace spanned by a test set. The resulting ratio, which ranges from 0 to 1, represents the fraction of conformational variance in the reference set that is captured by the test set. A key property of PCV is its asymmetry, which enables comparative evaluation between methods. Method B is considered superior to Method A when $\text{PCV}(A,B)$ exceeds $\text{PCV}(B,A)$. This indicates that the conformational space of Method B captures most of the variance present in Method A, while Method A fails to explore the conformations unique to Method B. Such a relationship indirectly demonstrates that Method B provides a more comprehensive and diverse characterization of the conformational landscape.
Method IgG-1D IgG-RL
GT 0.8091 0.8177 0.8258 0.7009 0.6294 0.7320
Table 1: PCV result on IgG datasets. ”–” indicates that the method fails to produce a valid result.
We first evaluated the performance of compared methods on the simulation IgG-1D and IgG-RL datasets, for which ground-truth (GT) conformations are available. For each method, an ensemble of 100 conformations was generated by applying K-means clustering to the latent vectors and decoding the cluster centroids into 3D volumes. As summarized in Table 1, GaussianEM yields a more diverse conformational ensemble than CryoDRGN and 3DFlex. In contrast, DynaMight fails to effectively resolve conformations in datasets with large-scale structural motions.
@@
Next, we extended our evaluation to the experimental EMPIAR datasets described in the main text. For each method, the conformational space was represented by 50 reconstructed conformations. As shown in Table 2, the comparison between GaussianEM and CryoDRGN (dark blue) indicates that GaussianEM recovers a broader and more diverse conformational space. For example, in the Supplementary Material section “T6SS effectors”, GaussianEM successfully captures the slight conformational motions of effectors that are not resolved by CryoDRGN. Similarly, the comparisons highlighted by light green and light blue demonstrate the superiority of CryoDRGN over DynaMight and of DynaMight over 3DFlex, respectively. Overall, these results establish a consistent performance ranking: GaussianEM >> CryoDRGN >> DynaMight >> 3DFlex. A similar ranking was observed across the other EMPIAR datasets.
@@
Finally, to further assess robustness and generalizability, we performed PCV comparisons on two datasets from the FI cryo-EM heterogeneity challenge, both of which exhibit complex conformational heterogeneity of thyroglobulin (Table 3). The first dataset is experimental and contains 33,742 particle images, while the second is a simulated dataset of the same size, generated by projecting 3,861 distinct conformations derived from molecular dynamics simulations. The pairwise PCV results consistently show that GaussianEM outperforms CryoDRGN in resolving conformational heterogeneity, with CryoDRGN in turn outperforming 3DFlex. Notably, DynaMight failed to produce plausible conformations on either dataset under default parameters.
FSC-based metrics on volume generation
Figure 7: a Sample-FSC results of compared methods on IgG datasets. b Cluster-FSC results of compared methods on IgG dataset.
CryoBench introduced an FSC-based metric, termed Sample FSC, to evaluate the reconstructed volumes. Sample FSC is computed by first applying K-means clustering to latent embeddings. The cluster centroids are then decoded into representative volumes. Because these unsupervised reconstructions do not directly correspond to GT volumes, the Area Under the Fourier Shell Correlation curve ($FSC_{AUC}$) is computed between each reconstructed volume and all GT volumes. The maximum value is retained as the score for each reconstruction. The distribution of these maximum scores across clusters reflects the overall generative performance of the model. Based on this idea, we introduce a new metric, termed Cluster FSC. Unlike Sample FSC, Cluster FSC clusters 2D projection images using their corresponding latent embeddings and subsequently reconstructs representative 3D volumes for each cluster by back-projection. While Sample FSC primarily evaluates the generative quality of the decoder, Cluster FSC provides a more rigorous assessment of the model’s classification and disentanglement capabilities. By reconstructing volumes directly from clustered projection images rather than from decoded latent variables, Cluster FSC metric reduces the impact of potential decoder-induced artifacts or biases. This design provides a more reliable assessment of whether conformational heterogeneity is correctly disentangled.
In this section, we use $K=20$ for all clustering tasks. Figure 7 shows the Sample FSC and Cluster FSC results for GaussianEM, 3DFlex, and CryoDRGN on the IgG-1D and IgG-RL datasets. As shown in Figure 7a, GaussianEM achieves higher reconstruction quality than CryoDRGN and 3DFlex, which is consistent with the visual results of decoded volumes (see Figure LABEL:fig:IgGd and Figure LABEL:fig:IgG_RLc in the Supplementary Material section “IgG simulation”). Figure 7b illustrates that GaussianEM outperforms CryoDRGN in conformational disentanglement and achieves performance comparable to 3DFlex. These results are consistent with the latent variable distributions exhibited in Figure 2b, Figure LABEL:fig:IgGc and Figure LABEL:fig:IgG_RL (presented in the Supplementary Material section “IgG simulation”).
@@ Discussion
In this study, we introduce GaussianEM, a Gaussian-based framework designed to intuitively capture both compositional and conformational heterogeneity in cryo-EM datasets. Most recent cryo-EM heterogeneity analysis methods typically rely on voxel-based representations of macromolecular structures. Such representations allow for the identification of compositional variations by comparing density between reconstructed volumes. However, they are often limited in describing continuous conformational motion. Tracking subtle structural changes using voxel grids usually requires manual interpretation and lacks smoothness. Additionally, this process sometimes introduces artifacts. In contrast, Gaussian-based models offer a more natural way to describe structural variability. In this formulation, changes in Gaussian density directly correspond to the presence or absence of components, while Gaussian displacements provide an intuitive visualization of continuous motion. This representation makes it easier to locate and interpret regions undergoing change. Moreover, the isotropic Gaussian function ensures consistent behavior in both two and three dimensions, linking 2D projections directly to 3D structural deformations. These properties form the foundation of GaussianEM in exploring molecular heterogeneity.
@@ Methods
We represent the protein density map with a group of learnable 3D Gaussians $\mathbb{G}^{3}=\{G^{3}_{i}\}_{i=1,\ldots,N}$ Kerbl et al. (2023). Each Gaussian $\mathbb{G}^{3}_{i}$ depicts a local Gaussian-shaped density distribution, i.e.,
@@
A good property of the Gaussian distribution is that its integral along one coordinate axis (usually $z$ row in the Cryo-EM domain) in 3D yields a 2D Gaussian distribution Zha et al. (2024); Bae et al. (2024). The $\hat{\mathbf{x}}\in\mathbb{R}^{2}$, $\hat{\mathbf{p}}_{i}\in\mathbb{R}^{2}$ and $\hat{\mathbf{\Sigma}}_{i}\in\mathbb{R}^{2\times 2}$ of projected Gaussians $\mathbb{G}^{2}_{i}$ are easily obtained by directly discarding the specific rows and columns of their corresponding 3D versions. The rendered particle projection $I_{k}\in\mathbb{R}^{H\times W}$ can be formulated as
@@
where $\mathcal{R}$ denotes the complete rendering process, $C_{k}$ denotes the CTF modulator, $P$ is the simple projection operator along the $z$ row, and $T$ transforms the 3D Gaussians given pose $\phi_{k}$. For computational efficiency, $T$ actually calculates the inverse matrix of the particle pose to align the viewing direction with the coordinate axis.
@@
Consistent with most existing work, the heterogeneous reconstruction network adopts an encoder–decoder architecture that maps raw particle images to their possible deformed volumes. Within the framework of 3D Gaussians, the deformation of flexible structures is characterized by the change of Gaussian parameters, whereas the invariant parameters correspond to the canonical density map obtained from homogeneous reconstruction. Our heterogeneous reconstruction network consists of three main modules, two encoders and one decoder, all implemented using multilayer perceptrons with Relu activation functions (see Figure LABEL:fig:network Supplementary Material section “Network architecture”). First, the raw particle images are processed by an image encoder made up of three hidden layers, which transforms them into $i$-dimension latent variables. In neural radiance field-based methods, position encoding is introduced to facilitate the network determining particle poses and learning high-frequency features. Similarly, we design a Gaussian encoder to assign a unique identifier to each Gaussian, helping to distinguish and aggregate different Gaussians. This encoder takes the density, scaling, position, and rotation parameters of each Gaussian as input and produces a $j$-dimension latent embedding. This per-Gaussian strategy has no limit on the number of Gaussians, allowing the incorporation of given atomic models and yielding more accurate estimations.
@@
Geometric Constraints. Overfitting can easily occur in the absence of well-designed regularization, particularly when dealing with high-noise experimental images. In such cases, Gaussians tend to aggregate, resulting in low-resolution structures that contribute uniformly across inputs. Considering the inherent continuity of protein motions, smoothness and local rigidity are essential properties to maintain. For this purpose, we impose two types of constraints on the estimated parameters: one for Gaussian embedding and the other just on the Gaussian displacement. For Gaussian embedding $\mathbf{j}_{n}$, given the average value of nearest point distance $d_{mean}$, we consider the neighboring Gaussians within a radius of $\mu_{1}=2.5\times d_{mean}$;
@@
where $\mathcal{N}(n,\mu)$ denotes the set of Guassians within distance $\mu$ from Gaussian $n$, and $d_{n,s}$ is the initial space distance between Gaussians $n$ and $s$. The weight $\omega_{n,s}$ controls the strength of constraints. The closer the two Gaussians are, the stronger the constraint is. This embedding constraint encourages adjacent Gaussians to undergo similar parameter changes, thereby preserving local consistency in both compositional and conformational heterogeneity. For the displacements of Gaussian, we first enforce the local rigidity of deformations by constraining the relative distance between adjacent Gaussians.